# Stiff Circuit System Modeling via Transformer

### v1

@@ 1. Introduction
In the modern circuit design industry, SPICE simulator is a cornerstone tool that enables accurate, early-stage validation of various designs. However, SPICE simulator relies heavily on physical equation-based models and iterative algorithms (Pratap et al., 2014), which makes it computationally expensive for the simulation of large scale circuits. Furthermore, physics models contain every detail of the design such as the numerical values of transistor sizes and capacitors, which is highly confidential. Design companies must share behavior models with downstream clients for integration and testing purposes, but if these models are leaked to outsiders, they can be easily reverse-engineered to reveal the original design. Therefore, a physics-based model will also pose risk to the protection of intellectual property (IP) and it is crucial to develop highly accurate black-box behavioral models as a replacement for physics-based models.
Among the alternatives to the physics-based model, Neural Network (NN) stands out as an ideal option, offering both generalizability and accuracy. Previous work  (Liu et al., 2004) uses NN to develop a black-box model for circuit transient simulation. In  (Luongvinh and Kwon, 2005; Chen et al., 2017), the authors used Recurrent Neural Network (RNN) to improve the model accuracy and proposed a way to enable the co-simulation of learned ML-based models and physics-based models. This makes ML-based models more practical to use. While  (Liu et al., 2004; Luongvinh and Kwon, 2005; Chen et al., 2017) are all discrete-time models,  (Xiong et al., 2021) leverages ODE-RNN  (Rubanova et al., 2019) to develop continuous-time circuit behavioral models to achieve the state-of-the-art performance. One remaining problem in  (Xiong et al., 2021) is that the model performance will be much worse when it is used to model a stiff circuit system. Stiffness refers to the situation when some parts of a system change very slowly, while other parts change very quickly. This is not a rare situation in circuit design. In fact, most of today’s circuits operate with input signals and output signals at very different frequencies, but unfortunately none of the previous ML-based circuit models can learn the stiff behavior effectively. Some previous works  (Visweswariah and Rohrer, 1989; Weng et al., 2012) tried to attack this problem via equation-based models, but they are hard to implement and not a universal solution to all types of circuits. Also, as we mentioned in the beginning, developing black-box model is more preferrable.
In this paper, we want to propose a new circuit behavioral model that can capture the stiff behavior of the circuit. The proposed model should be able to correctly predict the circuit output when it receives high frequency input, low frequency input, or a mixture of both. The model should have flexibility to model a wide range of circuits.
We propose to utilize the attention-mechanism to achieve this goal. In recent years, with the appearance of Transformer  (Vaswani et al., 2017), we have seen how attention mechanisms can be used to boost the model performance in generative tasks.  (Chien and Chen, 2021) extends ODE-RNN with the attention mechanism to prove that the attention mechanism can also enhance the model performance for time-series forecasting tasks. In  (Chen et al., 2023), the authors combined the Transformer with Neural ODE to enable continuous-time temporal data modeling. With the success of transformer, we believe that an attention-based circuit model can better express the behavior of stiff circuit system. We will leverage the architecture of the transformer proposed in  (Chen et al., 2023; Zhang and Yan, 2023) and explore different data representations methods for both high-frequency and low-frequency data, enabling the development of a more effective model. Since data samples are expensive to acquire, we will investigate active learning methods to reduce the amount of data needed for model training. In addition, we will derive the necessary conditions for learning a stable transformer-based model, which is a guaranteed feature in physics-based models. Since the model has built-in Neural ODE components, we can do mathematical analysis in the ODE equations to obtain the necessary conditions to be met for learning a stable model.
The proposed model will be evaluated based on its performance comparing with the previous state-of-the-art framework  (Xiong et al., 2021). We plan to test the performance of models on analog-to-digital-converter (ADC) circuit, which is one of the most common circuit in the integrated circuit industry. We may also consider adding other more complex circuits as test cases to provide a more comprehensive comparison. We will also benchmark the needed amount of data to prove this model is practical to use.
@@
Figure 1. Our model leverages Crossformer (Zhang and Yan, 2023) and KAN (Liu et al., 2024) for time-series forecasting. Specifically, Crossformer is used to incorporate input time-series information. Then, KAN is added to the decoder layer to project outputs to the desired forecasting dimensions.
This work demonstrates the effectiveness of using transformer architecture to model the stiff system, and the feasibility of using transformer-based surrogate model in circuit modeling problem. This model would greatly enhance the overall circuit design efficiency by speeding up simulations and reducing computational cost. In addition, the introduction of this model may inspire more development and use of ML-based surrogate model in the integrated circuit industry.
@@ 2. Related works
In recent years, with the appearance of Transformer  (Vaswani et al., 2017), we have seen how attention mechanisms can be used to boost the model performance in generative tasks.  (Chien and Chen, 2021) extends ODE-RNN with the attention mechanism to prove that the attention mechanism can also enhance the model performance for time-series prediction tasks. In  (Chen et al., 2023), the authors did not further extend the RNN structure. Instead, they leveraged the architecture of the Transformer and extended it to a continuous time domain.
@@ 3. Methodology
As shown in Fig. 1, we mainly leverage the structure of Crossformer  (Zhang and Yan, 2023), which is the state-of-the-art model for multivariate time-series forecasting. Then, we add Kolmogorov-Arnold Networks (KANs)  (Liu et al., 2024) layers to the last layer of the Crossformer decoder to project the hidden vectors to the dimension of the circuit model.
@@ 7. Conclusion
In this work, we proposed a new circuit behavioral model based on KANs and Crossformer, which outperforms the previous state-of-the-art model CTRNN for the stiff circuit modeling task. Using the same dataset, our model is able to provide a more accurate prediction with less training time. In the future, we may consider to add Fourier transform in feature embedding to further increase the model accuracy.

### v2

@@ 1. Introduction
In the modern circuit design industry, the SPICE simulator is a cornerstone tool that enables accurate, early-stage validation of various designs. However, SPICE relies heavily on physics-based models and iterative numerical algorithms (Pratap et al., 2014), making it computationally expensive for large-scale circuit simulations. In addition, these models encode detailed design parameters such as transistor sizes and capacitances, which are highly confidential. When behavioral models are shared with downstream clients for integration and testing, such detailed information can potentially be reverse-engineered, posing risks to intellectual property (IP) protection. These limitations motivate the development of accurate black-box behavioral models that can approximate circuit responses without exposing internal design details.
Neural networks (NNs) have emerged as a promising alternative due to their strong approximation capability and flexibility. Prior work (Liu et al., 2004) applies NNs to circuit transient simulation, while (Luongvinh and Kwon, 2005; Chen et al., 2017) further improves accuracy using recurrent neural networks (RNNs) and enables co-simulation with physics-based models. Although these approaches demonstrate improved practicality, they are primarily designed for discrete-time modeling. To address this limitation, (Xiong et al., 2021) leverages ODE-RNN (Rubanova et al., 2019) to develop continuous-time circuit behavioral models, achieving state-of-the-art performance.
Despite these advances, modeling stiff circuit systems remains a significant challenge. Stiffness refers to systems in which certain components evolve slowly while others change rapidly. This phenomenon is common in circuit design, where high-frequency signals (e.g., clock edges) coexist with low-frequency analog dynamics. Existing ML-based models often struggle to capture such heterogeneous behavior, leading to degraded performance when applied to stiff systems. While prior works (Visweswariah and Rohrer, 1989; Weng et al., 2012) attempt to address stiffness using equation-based approaches, they are often difficult to implement and lack generality across different circuit types. Therefore, developing a robust black-box model that can effectively capture stiff dynamics remains an open problem.
From another perspective, stiff circuit behavior can be interpreted as inherently multi-modal, where different temporal components—such as fast switching transients and slow-varying signal envelopes—exhibit distinct characteristics. In this view, circuit responses are governed by multiple interacting modalities rather than a single homogeneous process. Such heterogeneity makes it difficult for conventional models to learn a unified representation. Recent advances in representation learning (Du et al., 2026; Jing et al., 2026; Li et al., 2025e; Liu et al., 2024a) suggest that capturing interactions across different modalities, while preserving their unique characteristics, can significantly improve modeling fidelity. This perspective provides a useful lens for understanding the complexity of stiff circuit systems.
In this paper, we propose a new circuit behavioral model that is capable of capturing stiff dynamics across a wide range of input conditions. The model is designed to accurately predict circuit outputs under high-frequency inputs, low-frequency inputs, or mixtures of both, while maintaining flexibility across different circuit types.
To achieve this goal, we leverage attention-based architectures. With the emergence of Transformer models (Vaswani et al., 2017), attention mechanisms have demonstrated strong performance in sequence modeling and time-series prediction. Prior work (Chien and Chen, 2021) extends ODE-RNN with attention to improve time-series forecasting, while (Chen et al., 2023) combines Transformer architectures with Neural ODEs for continuous-time modeling. Motivated by these successes, we hypothesize that attention mechanisms are well-suited for capturing the multi-scale and multi-modal nature of stiff circuit dynamics. Specifically, we build upon the Transformer architectures proposed in (Chen et al., 2023; Zhang and Yan, 2023) and explore data representations tailored for both high-frequency and low-frequency signals.
@@
Figure 1. Our model leverages Crossformer (Zhang and Yan, 2023) and KAN (Liu et al., 2024b) for time-series forecasting. Specifically, Crossformer is used to incorporate input time-series information. Then, KAN is added to the decoder layer to project outputs to the desired forecasting dimensions.
This work demonstrates the effectiveness of transformer-based architectures for modeling stiff circuit systems and highlights the potential of attention-based surrogate models in circuit design. By improving simulation efficiency and reducing computational cost, the proposed approach can significantly enhance circuit design workflows. Moreover, this work provides a new perspective on treating circuit dynamics as multi-modal temporal processes, which may inspire further research on machine learning-based surrogate modeling in the integrated circuit industry.
@@ 2. Related works
In recent years, with the emergence of Transformer models (Vaswani et al., 2017), attention mechanisms have demonstrated strong capability in improving performance on time-series tasks (Liu et al., 2025, 2026; Ma et al., 2025). For example, (Chien and Chen, 2021) incorporates attention into the ODE-RNN framework, showing that attention can enhance prediction accuracy in time-series modeling. Rather than extending recurrent architectures,  (Chen et al., 2023) directly adopts the Transformer paradigm and extends it to the continuous-time domain, enabling more effective modeling of temporal dynamics. In addition, other approaches from the time-series community have explored alternative generative frameworks such as GANs (Zhang et al., 2026). However, these methods have not been widely investigated in the context of circuit behavioral modeling.
@@ 3. Methodology
As shown in Fig. 1, we mainly leverage the structure of Crossformer  (Zhang and Yan, 2023), which is the state-of-the-art model for multivariate time-series forecasting. Then, we add Kolmogorov-Arnold Networks (KANs)  (Liu et al., 2024b) layers to the last layer of the Crossformer decoder to project the hidden vectors to the dimension of the circuit model.
@@ 7. Conclusion
In this work, we proposed a transformer-based circuit behavioral model that integrates KANs and Crossformer, achieving superior performance over the prior state-of-the-art CTRNN model for stiff circuit modeling. Under the same dataset, our approach delivers higher prediction accuracy while requiring less training time.
Looking forward, one promising direction is to incorporate more domain knowledge (Kang et al., 2020; Hu et al., 2026; Wu et al., 2022; Yan et al., 2023, 2025), such as Fourier-based representations  (Wu et al., 2025) in the feature embedding stage, to further enhance model performance. Another potential avenue is to introduce adaptive refinement mechanisms during training and inference, inspired by recent modular self-improving and planning-based frameworks in machine learning (Yang et al., 2026b; Li et al., 2025f, 2026; Yang et al., 2026a; Jia and Li, 2024).
Finally, we note that, unlike many other popular research areas (Cai et al., 2025; Wang et al., 2025; Guo et al., 2025; Li et al., 2025b, a), there is currently no publicly accessible dataset specifically designed for stiff circuit modeling. Establishing standardized public datasets would greatly facilitate benchmarking and promote further research in this domain. Data augmentation (Li et al., 2025d; Wang and Li, 2026; Li et al., 2025c) or active learning (Shen et al., 2021) technique might be used as a temporary solution for the lack of data.